OpenStudy (anonymous):

At the beginning of this month, the balance of Agatha's checking account was $782.39. So far this month, she has received a paycheck via direct deposit of $932.48, been charged a monthly service fee from her bank of $20.00, used a debit card linked to her account to make a purchase of $36.82, written a check for $155.03 that has already been deposited, and deposited a check written to her for $79.13. What is the current balance of Agatha's checking account? A. $1602.15 B. $1582.15 C. $799.76 D. $1733.95 @Luigi0210

OpenStudy (debbieg):

---The balance of Agatha's checking account was $782.39. That's the amount she starts with in her account --- received a paycheck via direct deposit of $932.48 That gets added to her balance. --- been charged a monthly service fee from her bank of $20.00 That's a FEE CHARGED to her, so what do you think - added or subtracted from her balance? --- used a debit card linked to her account to make a purchase of $36.82 Purchase via debit card, so that is "debited" (deducted) from her account balance. --- written a check for $155.03 that has already been deposited She wrote the check from her account, AND it was deposited, so her bank has processed that check and taken it OUT of her account, so it's SUBTRACTED from her balance. --- and deposited a check written to her for $79.13. Another deposit - that's an ADDITION to her balance.

OpenStudy (anonymous):

im confused @DebbieG

OpenStudy (luigi0210):

Oh darn, I'm late to Debbie's lecture.

OpenStudy (debbieg):

Why? what did you not understand above? She starts with a balance in the account. some things are added to that, some are deducted from that. You just need to add and subtract... tell me which item you are confused on, I'll help you.

OpenStudy (debbieg):

lol 1 demerit for you @Luigi0210 :P

OpenStudy (anonymous):

your just repeating the question

OpenStudy (debbieg):

No, I'm not. I made remarks about each item, to help you think about what gets ADDED and what gets SUBTRACTED. Did you read everything that I wrote?

OpenStudy (anonymous):

yes ill reread

OpenStudy (anonymous):

i got 1423.89

OpenStudy (debbieg):

Lets do the first one: ---The balance of Agatha's checking account was $782.39. That's the amount she starts with in her account --- received a paycheck via direct deposit of $932.48 That gets added to her balance. So at this point her balance started with 782.39, and then the deposit was ADDED to her balance, so now she has: 782.39 + 932.48 = 1714.87 Now just keep going through the transactions... add the things that are additions to the account, and subtract the things that are coming OUT of the account.

OpenStudy (debbieg):

ALMOST. you have ONE transaction wrong. "--- and deposited a check written to her for $79.13. Another deposit - that's an ADDITION to her balance." You subtracted that, instead of adding it.... deposits to her account are ADDITIONS.

OpenStudy (anonymous):

i got 1794

OpenStudy (anonymous):

@DebbieG

OpenStudy (debbieg):

No. Can you show your work? Just show me each set of addition/subtraction. Something is going wrong somewhere.... you were only off by that one transaction the first time, but now you're off by more than that.

OpenStudy (anonymous):

i did 1714.87+79.13

OpenStudy (anonymous):

@DebbieG

OpenStudy (debbieg):

1714.87 was from MY example of just the FIRST transaction. There were a bunch of transactions between that and the error of 79.13. show your work from the beginning. each transaction.

OpenStudy (anonymous):

omg your confusing me

OpenStudy (anonymous):

can you help with numbers instead of words

OpenStudy (debbieg):

look @divagirl421 , I won't do this problem FOR you. This is basic math, and understanding additions and subtractions to a checking account is a basic life skill. You added and subtracted some of these numbers to get your answer - I'm just asking YOU to show ME the numbers, and then I can see where you went wrong so that you can LEARN from it. Me doing the problem FOR you won't teach you anything.

OpenStudy (anonymous):

im not telling you to do that im just confused

OpenStudy (debbieg):

I don't know how I can help you any more, without doing the problem. I told you what to do with EACH number that is in the problem. I did the first transaction for you to demonstrate. I told you exactly what error you made with your first answer. I know you are frustrated, but I can't help you any more if you won't simply show me the math that you did to get your answer.

OpenStudy (anonymous):

782.39.+$932.48=1714.87 1714.87-20=1694.87 1684.87-36.82 1658.05-155.03=1503.02 1503.02+79.13=1582.15 idk if i did it right

OpenStudy (debbieg):

THAT'S IT!! That's all you had to do...just the correct +'s and -'s!!!

OpenStudy (debbieg):

\(\begin{matrix}\Huge \color{gold}{★★★★★★★★★★}\\ \Huge \color{red}{\text{Great job!}} \\ \Huge \color{red}{\text{Gold stars for you!}}\\ \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix} \color{purple}{\overbrace{\left(\begin{matrix}\color{blue}{\widehat\odot \widehat\odot} \\\small\bigtriangleup \\ \Huge \color{red}\smile\end{matrix}\right)}} \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix}\end{matrix}\)

OpenStudy (anonymous):

oooo yayyyyy thank you sooo much:)

OpenStudy (debbieg):

you're welcome. :) now wasn't that worth the effort? :)

OpenStudy (anonymous):

yes:)